OpenStudy (halfdeafdarling):

I need major help figuring out how to explain this in sentences, I know the answer is (2,2) but the way I do it is with addition and elimination. Can someone explain to me how to do it with multiplication first? Explain, in complete sentences, how you would use the elimination method to solve the following system of equations and provide the solution to the system. 5x - 3y = 4 3x - 5y = -4

OpenStudy (danjs):

If you have a system of linear equations like given above, you can do 3 things to change it. 1-rearrange order of equations 2-multiply an equation by a number 3-add one equation to another equation. This elimination method involves adding the two equations together, and having the right numbers so that either the x or y will go to zero and fall out. 5x - 3y = 4 3x - 5y = -4

OpenStudy (danjs):

say you want to eliminate the y variable. Multiply equation 1 by 5 , and multiply equation 2 by -3. Then you add the two equations together and the y variable will go to zero. 25x - 15y = 20 -9x + 15y = 12 ----------------+ 16x + 0y = 32 that gives you 16x = 32 or x=2, and using x=2 in any equation, y=2